# A Novel Garment Transfer Method Supervised by Distilled Knowledge of Virtual Try-on Model

### v1

@@ 3.2.3 Student Transfer Parsing Reasoning Model
Typically, when a teacher teaches a student how to solve a problem, they first introduce the basic knowledge and mechanism, followed by a solution procedure from implicit to explicit, and provide feedback to supervise the student’s learning. If the student aims to become proficient at solving this type of problem, they must practice more through self-study. Drawing inspiration from this, we delicately design the knowledge distillation between parsing reasoning models as three phases: feature knowledge teaching, response knowledge teaching, and self-study phase.
@@
We adopt the same backbone in ${{{\cal S}_{p}}}$ as in ${{{\cal T}_{p}}}$. During the feature and response knowledge teaching phase, ${{{\cal S}_{p}}}$ is conditioned on ${{\cal M}_{B}^{hc}}$, ${{{\cal P}_{A}}}$, and ${{\cal M}_{A}^{hr}}$. Due to the unaligned pose of person ${B}$ in ${{\cal M}_{B}^{hc}}$ compared to ${{\cal M}_{B}^{c}}$, transfer parsing reasoning poses a greater challenge compared to try-on parsing reasoning. As Fig. 3 shows, the encoder in the backbone is responsible for feature extraction and position awareness. However, since the input between the teacher and student models differs, their learned knowledge in the encoder may not be identical. On the other hand, the decoder in the backbone is responsible for entangling semantic features and reconstructing the parsing structure, and the teacher and student models share the same objective for it. Therefore, we propose to distill the decoder’s knowledge from ${{{\cal T}_{p}}}$ to supervise the training of ${{{\cal S}_{p}}}$.
@@
Specifically, we focus on the layers ’hx6’, ’hx5d’, ’hx4d’, ’hx3d’, ’hx2d’, and ’hx1d’  [^1] These signs are derived from the official U2-Net code (https://github.com/NathanUA/U-2-Net). in the decoder, which represent a sequence of outputs ranging from coarse to fine. As Fig. 4 shows, it indicates that the global pose and overall shape are learned at the early layers, ’hx6’ and ’hx5d’, which are considered basic knowledge for reasoning. Subsequently, the model starts to re-establish human structure and distinguish categories at the intermediate layers, ’hx4d’, ’hx3d’, and ’hx2d’, with particular attention to the boundary and overall shape of the garment and upper body skin, which represent the entanglement of the garment and pose features. Finally, at layer ’hx1d’, our models refine the shape of garment-agnostic categories by fusing low-level shape information from skip connections. During the first stage, we only distilled knowledge from layers ’hx6’, ’hx5d’, which indirectly supervises ${{{\cal S}_{p}}}$ at the extraction of hidden knowledge of garment style and target pose. During the second stage, we distilled knowledge from all layers, which supervises ${{{\cal S}_{p}}}$ at the entanglement of semantic features and reconstruction of parsing spatial relationships.
During the response knowledge teaching phase, we distill soft knowledge in the form of a heat map instead of using one-hot coding. Since the soft target represented by ${{{\cal M}_{TAB}}}$ is not equivalent to the ground truth, ${{{\cal S}_{p}}}$ is able to acquire hard knowledge through self-study after the teaching phase, thereby improving its overall reasoning capacity and facilitating the correction of teaching errors. Specifically, we split the garment category mask ${{\cal M}_{SAB}^{hc}}$ from ${{{\cal M}_{SAB}}}$, re-transfer it onto person ${B}$, and supervise the reasoning result by the ground truth ${{{\cal M}_{B}}}$. It is described as ${{{\cal M}_{SB}}={{\cal S}_{p}}\left({{{\cal P}_{B}},{\cal M}_{B}^{hr},{\cal M}_{SAB}^{hc}}\right)}$ in the formula.
@@ 3.3.1 Garment Warping Insight
After reasoning the transfer parsing ${{{\cal M}_{SAB}}}$, we exploit it as a shape prior and guide the garment warping, which enables the garment ${{\cal I}_{B}^{hc}}$ of person ${B}$ to conform to the target shape ${{\cal M}_{SAB}^{hc}}$. To achieve this, there are differentiable warping tools available such as STN (Jaderberg et al., 2015), TPS (Bookstein, 1989), and optical flow (Dosovitskiy et al., 2015), which learn affine transformation, coarse-grained grid movement, and pixel displacement, respectively, with varying degrees of freedom. However, for garment transfer, the warping needs to match the garment shape with the target shape in another pose while preserving the garment’s style and texture. Thus, the fine-grained optical flow is the most suitable tool for this task.
Although we derive the target shape for garment warping, the ground truth of garment warping ${{\cal I}_{AB}^{hc}}$ remains unavailable. Supervising the training of the flow warping model for garment transfer with only shape information may lead to distortion and loss of garment semantic information. Therefore, adequate supervision with content information is crucial for estimating high-DOF optical flow. As paired data ${\left({{\cal I}_{A}^{c},{\cal I}_{A}^{hc}}\right)}$ is available, we can train the flow warping model for virtual try-on in a self-supervised manner, and we supervise the flow warping model for garment transfer with both shape and content information by distilling knowledge from the flow warping model for virtual try-on.
@@
Specifically, as Fig. 5 shows, we leverage the parser-based version of FSV (He et al., 2022) as the teacher flow warping model ${{{\cal T}_{w}}}$, which is capable of estimating global style-based flow and refining flow locally. During its training phase, we represent the human and garment utilizing ${{\cal M}_{A}^{hc}}$ and ${\left({{\cal M}_{A}^{c},{\cal I}_{A}^{c}}\right)}$, and exploit the ground truth ${{\cal I}_{A}^{hc}}$ to supervise the warping result ~${\tilde{\cal I}_{A}^{hc}}$. Once ${{{\cal T}_{w}}}$ is trained, we exploit it to warp garment-in-shop ${B}$ onto person ${A}$. i.e. ~${\tilde{\cal I}_{TAB}^{hc}={{\cal T}_{w}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{B}^{c},{\cal I}_{B}^{c}}\right)}$. With the aid of the supervision information ${{\cal M}_{SAB}^{hc},\tilde{\cal I}_{TAB}^{hc}}$ at the shape and content levels, we delicately devise a student flow warping model for garment transfer, inspired by the progressive flow estimation technique in (Ge et al., 2021b), which has the initial aligning phase and the flow warping phase.
@@ 3.3.2 Initial Aligning Phase
To address the garment warping problem effectively, we divide it into two subtasks, namely position mapping and shape adjustment. The former task modifies the global position of the garment by scaling, translation, and rotation, while the latter adjusts the local shape to match the target shape precisely using non-rigid warping. For virtual try-on applications, the garment-in-shop usually has a standard pose and a high pixel ratio, as shown in Fig. 6(a), which enables the flow warping model to perform both position mapping and shape adjustment tasks. However, for garment transfer, as shown in Fig. 6(b), the diversity of cropping and pose in the person image may cause the transferred garment to deviate significantly from the target shape in terms of both position and shape. To address this issue, we propose to utilize STN to perform position mapping independently, as shown in Fig. 6(c), which reduces the learning burden of the subsequent flow warping task. Furthermore, due to the locality of CNN, the flow warping model is not able to capture the long-range shape dependence, which may lead to distortion during precise shape matching, as shown in Fig. 6(d). However, the initial aligning phase can mitigate this issue by applying the global affine transformation.
@@ 3.3.3 Flow Warping Phase
After the initial aligning, the flow warping phase learns correspondence from coarse to fine levels and estimates a refined flow to accurately warp garment ¯${\bar{\cal I}_{B}^{hc}}$ to target shape ${{\cal M}_{SAB}^{hc}}$ precisely. To achieve this, the shared encoder ${{E_{\gamma}}}$ extracts the pyramid feature maps ${\left\{{\gamma_{1}^{i}}\right\}_{i=1}^{i=5}}$ and ${\left\{{\gamma_{2}^{i}}\right\}_{i=1}^{i=5}}$ from weak-sematic shape representation ${\left\{{{\cal M}_{SAB}^{hc},{{\cal P}_{A}}}\right\}}$ and {¯${\left\{{\bar{\cal M}_{B}^{hc},{{\bar{\cal P}}_{B}}}\right\}}$ at five downsampling layers. The encoder ${{E_{\delta}}}$ has the same structure as ${{E_{\gamma}}}$ but does not share weights, which extracts the pyramid feature maps ${\left\{{\delta_{1}^{i}}\right\}_{i=1}^{i=5}}$ from ¯${\bar{\cal I}_{B}^{hc}}$. Consequently, we obtain the paired shape feature maps ${\left\{{\gamma_{1}^{i},\gamma_{2}^{i}}\right\}}$ and the paired content-shape feature maps ${\left\{{\gamma_{1}^{i},\delta_{1}^{i}}\right\}}$, which serve as the condition of the subsequent shape-content-aware warping module (SCAW).
Five SCAWs learn the correspondence between feature maps at the shape level and content level. Each SCAW consists of a shape-aware stage and a content-aware stage. For convenience, we index the downsampling layers of encoders from ${i=1}$ to ${i=5}$ and index SCAW from ${i=5}$ to ${i=1}$. Consequently, the ${i}$-th SCAW has the same scale as the ${i}$-th downsampling layer of encoders and is conditioned on the last flow ${f_{S}^{i+1}}$, ${\gamma_{1}^{i}}$, ${\gamma_{2}^{i}}$, and ${\delta_{1}^{i}}$. At the shape-aware stage, we first warp ${\gamma_{2}^{i}}$ to ${\dot{\gamma}_{2}^{i}}$ via ${f_{S}^{i+1}}$ and calculate the shape cost volume ${c_{1}^{i}}$ using ${\dot{\gamma}_{2}^{i}}$ and ${\gamma_{1}^{i}}$ as described in (Ilg et al., 2017). We then estimate the shape flow residual ${\nabla\dot{f}_{S}^{i}}$ using four Conv2d+LeakyRelu combinations and add it to ${f_{S}^{i}}$ to obtain the shape-aware flow ${\dot{f}_{S}^{i}}$. At the content-aware stage, we similarly warp ${\delta_{1}^{i}}$ to ${\dot{\delta}_{1}^{i}}$ via ${\dot{f}_{S}^{i}}$ and calculate the content cost volume ${c_{2}^{i}}$ using ${\dot{\delta}_{1}^{i}}$ and ${\gamma_{1}^{i}}$. The content flow residual ${\nabla\ddot{f}_{S}^{i}}$ is estimated using another four Conv2d+LeakyRelu combinations and added to ${\dot{f}_{S}^{i}}$ to obtain the upsampled output ${f_{S}^{i}}$. Moreover, the first SCAW module initializes the flow by calculating ${c_{1}^{i}}$ and ${c_{2}^{i}}$ from ${\left\{{\gamma_{1}^{i},\gamma_{2}^{i}}\right\}}$ and ${\left\{{\gamma_{1}^{i},\delta_{1}^{i}}\right\}}$.
@@ 3.3.4 Objective Function
${{\ell_{wi}}}$ punishes the global shape matching in the initial aligning phase, which consists of two items as (5). The former item punishes the shape error between ¯${\bar{\cal M}_{B}^{hc}}$ and ${{\cal M}_{SAB}^{hc}}$, whereas the latter encourages the garment pixel ratio of ¯${\bar{\cal M}_{B}^{hc}}$ to be similar to ${{\cal M}_{SAB}^{hc}}$, where ${H}$ and ${W}$ represent the height and width of masks; ${{\lambda_{4}}}$ and ${{\lambda_{5}}}$ are weights.
@@
${{\ell_{wr}}}$ regularizes the flow displacement and encourages flow coordination, which consists of two items as (7). We exploit both the one-order constraint in (He et al., 2022) and the second-order constraint in (Ge et al., 2021b) to be ${{\ell_{wr1}}}$ and ${{\ell_{wr2}}}$. ${{\lambda_{9}}}$ and ${{\lambda_{10}}}$ are weights.
@@ 3.4 Arm Regrowth
During garment transfer, the shape of the upper body skin may also undergo changes. Therefore, it is crucial to determine an appropriate upper body skin to achieve transfer realism after garment warping. Based on a comparison of the upper body skin mask ${{\cal M}_{A}^{ha}}$ of person ${A}$ with the target upper body skin mask ${{\cal M}_{SAB}^{ha}}$, this task is divided into two cases: covering existing skin or exposing new skin. As shown in Fig. 7, the former scenario involves simply multiplying ${{\cal M}_{SAB}^{ha}}$ with ${{{\cal I}_{A}}}$, However, the latter scenario presents a difficult task that cannot be resolved through pixel-level operations. Consequently, it is imperative to design a model capable of reasoning about new upper body skin, which has been designated as ’arm regrowth’.
Figure 7: Two cases in training ${{G_{r}}}$.
The arm regrowth model ${{G_{r}}}$ is trained using an innovative training approach. In the training phase, both scenarios involving exposing new skin are achieved by adjusting the input. For the former scenario, the source and target masks are swapped, where ${{\cal M}_{A}^{ha}}$ and ${{\cal M}_{SAB}^{ha}\odot{\cal I}_{A}^{ha}}$ are taken as the target mask and source content. They are then fed into ResUnet (Zhang et al., 2018) to predict ~${\tilde{\cal I}_{A}^{ha}}$, guided by the ground truth ${{\cal M}_{A}^{ha}\odot{\cal I}_{A}}$. For the latter scenario, ${{\cal M}_{A}^{ha}}$ is randomly cropped to ˙${\dot{\cal M}_{A}^{ha}}$, with the half-plane shape removed on the horizontal or vertical axis. Using ${{\cal M}_{A}^{ha}}$ and ˙${\dot{\cal M}_{A}^{ha}\odot{{\cal I}_{A}}}$ as the target mask and source content, they are fed into the same ResUnet to predict ~${\tilde{\cal I}_{A}^{ha}}$. In the testing phase, ${{\cal I}_{A}^{ha},{\cal M}_{SAB}^{ha}}$ are inputted into the pre-trained ResUnet to yield ~${\tilde{\cal I}_{SAB}^{ha}}$, which is denoted as ~${\tilde{\cal I}_{SAB}^{ha}={G_{r}}\left({{\cal I}_{A}^{ha},{\cal M}_{SAB}^{ha}}\right)}$ in the formula.

### v2

@@ 3.2.3 Student Transfer Parsing Reasoning Model
Typically, when a teacher teaches a student how to solve a problem, they first introduce the basic knowledge and mechanism, followed by a solution procedure from implicit to explicit, and provide feedback to supervise the student’s learning. If the student aims to be skilled in this type of problem, they need to do more exercises in self-study. Drawing inspiration from this, we delicately design the knowledge distillation between parsing reasoning models as three phases: feature knowledge teaching, response knowledge teaching, and self-study phase.
@@
We adopt the same backbone in ${{{\cal S}_{p}}}$ as in ${{{\cal T}_{p}}}$. During the feature and response knowledge teaching phase, ${{{\cal S}_{p}}}$ is conditioned on ${{\cal M}_{B}^{hc}}$, ${{{\cal P}_{A}}}$, and ${{\cal M}_{A}^{hr}}$. Due to the unaligned pose of person ${B}$ in ${{\cal M}_{B}^{hc}}$ compared to ${{\cal M}_{B}^{c}}$, transfer parsing reasoning poses a greater challenge compared to try-on parsing reasoning. As Fig. 3 shows, the encoder in the backbone is responsible for feature extraction and position awareness. However, with the input difference between the teacher and student models, their learned knowledge in the encoder may not be identical. On the other hand, the decoder in the backbone is responsible for entangling semantic features and reconstructing the parsing structure, the teacher and student models share the same objective for it.. Therefore, we propose to distill the decoder’s knowledge from ${{{\cal T}_{p}}}$ to supervise the training of ${{{\cal S}_{p}}}$.
@@
Specifically, we focus on the layers ’hx6’, ’hx5d’, ’hx4d’, ’hx3d’, ’hx2d’, and ’hx1d’  [^1] These signs are derived from the official U2-Net code (https://github.com/NathanUA/U-2-Net). in the decoder, which output feature maps ranging from coarse to fine. As Fig. 4 shows, it indicates that the global pose and overall shape are learned at the early layers, ’hx6’ and ’hx5d’, which are considered basic knowledge for reasoning. Subsequently, the model starts to re-establish human structure and distinguish categories at the intermediate layers, ’hx4d’, ’hx3d’, and ’hx2d’, with particular attention to the boundary and overall shape of the garment and upper body skin, which represent the entanglement of the garment and pose features. Finally, at layer ’hx1d’, our models refine the shape of garment-agnostic categories by fusing low-level shape information from skip connections. During the first stage, we only distilled knowledge from layers ’hx6’, ’hx5d’, which indirectly supervises ${{{\cal S}_{p}}}$ to extract hidden knowledge of garment style and target pose. During the second stage, we distilled knowledge from all layers, which supervises ${{{\cal S}_{p}}}$ at the entanglement of semantic features and reconstruction of parsing spatial relationships.
During the response knowledge teaching phase, we distill soft knowledge in the form of a heat map instead of one-hot coding. Since the soft target represented by ${{{\cal M}_{TAB}}}$ is not equivalent to the ground truth, ${{{\cal S}_{p}}}$ is able to acquire hard knowledge through self-study after the teaching phase, thereby improving its overall reasoning capacity and facilitating the correction of teaching errors. Specifically, we split the garment category mask ${{\cal M}_{SAB}^{hc}}$ from ${{{\cal M}_{SAB}}}$, re-transfer it onto person ${B}$, and supervise the reasoning result by the ground truth ${{{\cal M}_{B}}}$. It is described as ${{{\cal M}_{SB}}={{\cal S}_{p}}\left({{{\cal P}_{B}},{\cal M}_{B}^{hr},{\cal M}_{SAB}^{hc}}\right)}$ in the formula.
@@ 3.3.1 Garment Warping Insight
After reasoning the transfer parsing ${{{\cal M}_{SAB}}}$, we exploit it as a shape prior and guide the garment warping, which enables the garment ${{\cal I}_{B}^{hc}}$ of person ${B}$ to conform to the target shape ${{\cal M}_{SAB}^{hc}}$. To achieve this, there are three dominant differentiable warping tools: STN [23], TPS [24], and optical flow [25, 26]. They learn affine transformation, coarse-grained grid movement, and pixel displacement, respectively, with varying degrees of freedom. However, for garment transfer, the warping tool is required to adapt the shapes between the garment image and the target mask while preserving the style and texture of the garment. Thus, the fine-grained optical flow is the most suitable tool for this task.
Although we derive the target shape for garment warping, the ground truth of garment warping ${{\cal I}_{AB}^{hc}}$ remains unavailable. Supervising the flow warping model for garment transfer only with shape information may lead to distortion and information loss of the garment. Therefore, the content information is crucial to supervising high-DOF optical flow estimation. As paired data ${\left({{\cal I}_{A}^{c},{\cal I}_{A}^{hc}}\right)}$ is available, we can train the flow warping model for virtual try-on in a self-supervised manner, and we supervise the flow warping model for garment transfer with both shape and content information by distilling knowledge from the flow warping model for virtual try-on.
@@
Specifically, as Fig. 5 shows, we leverage the parser-based version of FSV [11] as the teacher flow warping model ${{{\cal T}_{w}}}$, which is capable of estimating global style-based flow and refining flow locally. During its training phase, we represent the human and garment with ${{\cal M}_{A}^{hc}}$ and ${\left({{\cal M}_{A}^{c},{\cal I}_{A}^{c}}\right)}$, and exploit the ground truth ${{\cal I}_{A}^{hc}}$ to supervise the warping result ~${\tilde{\cal I}_{A}^{hc}}$. Once ${{{\cal T}_{w}}}$ is trained, we exploit it to warp garment-in-shop ${B}$ onto person ${A}$. i.e. ~${\tilde{\cal I}_{TAB}^{hc}={{\cal T}_{w}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{B}^{c},{\cal I}_{B}^{c}}\right)}$. With the supervision information ${{\cal M}_{SAB}^{hc},\tilde{\cal I}_{TAB}^{hc}}$ at the shape and content levels, we delicately devise a student flow warping model for garment transfer, inspired by the progressive flow estimation technique in [9], which has the initial aligning phase and the flow warping phase.
@@ 3.3.2 Initial Aligning Phase
We divide the garment warping problem into two subtasks, namely position mapping and shape adjustment. The former task modifies the global position of the garment by scaling, translation, and rotation, while the latter task adjusts the local shape to match the target shape precisely by non-rigid warping. For virtual try-on applications, the garment-in-shop usually has a standard pose and a high pixel ratio, as shown in Fig. 6(a), which enables the flow warping model to perform both position mapping and shape adjustment tasks. However, for garment transfer, as shown in Fig. 6(b), the diversity of cropping and pose of the personal image causes dramatic variation in position and shape. To address this issue, we propose to utilize STN to perform position mapping independently, as shown in Fig. 6(c), which reduces the learning burden of the subsequent flow warping task. Furthermore, due to the locality of CNN, the flow warping model is not able to capture the long-range shape dependence, which may lead to distortion during precise shape matching, as shown in Fig. 6(d). However, the initial aligning phase can mitigate this issue by applying the global affine transformation.
@@ 3.3.3 Flow Warping Phase
After the initial aligning, the flow warping phase learns correspondence from coarse to fine levels and estimates a refined flow to warp garment ¯${\bar{\cal I}_{B}^{hc}}$ to target shape ${{\cal M}_{SAB}^{hc}}$ precisely. To achieve this, the shared encoder ${{E_{\gamma}}}$ extracts the pyramid feature maps ${\left\{{\gamma_{1}^{i}}\right\}_{i=1}^{i=5}}$ and ${\left\{{\gamma_{2}^{i}}\right\}_{i=1}^{i=5}}$ from weak-sematic shape representation ${\left\{{{\cal M}_{SAB}^{hc},{{\cal P}_{A}}}\right\}}$ and {¯${\left\{{\bar{\cal M}_{B}^{hc},{{\bar{\cal P}}_{B}}}\right\}}$ at five downsampling layers. The encoder ${{E_{\delta}}}$ has the same structure as ${{E_{\gamma}}}$ but does not share weights, which extracts the pyramid feature maps ${\left\{{\delta_{1}^{i}}\right\}_{i=1}^{i=5}}$ from ¯${\bar{\cal I}_{B}^{hc}}$. Consequently, we obtain the paired shape feature maps ${\left\{{\gamma_{1}^{i},\gamma_{2}^{i}}\right\}}$ and the paired content-shape feature maps ${\left\{{\gamma_{1}^{i},\delta_{1}^{i}}\right\}}$, which serve as the condition of the subsequent shape-content-aware warping module (SCAW).
Five SCAWs learn the correspondence between feature maps at the shape level and content level. Each SCAW consists of a shape-aware stage and a content-aware stage. For convenience, we index the downsampling layers of encoders from ${i=1}$ to ${i=5}$ and index SCAW from ${i=5}$ to ${i=1}$. Consequently, the ${i}$-th SCAW has the same scale as the ${i}$-th downsampling layer of encoders and is conditioned on the last flow ${f_{S}^{i+1}}$, ${\gamma_{1}^{i}}$, ${\gamma_{2}^{i}}$, and ${\delta_{1}^{i}}$. At the shape-aware stage, we first warp ${\gamma_{2}^{i}}$ to ${\dot{\gamma}_{2}^{i}}$ via ${f_{S}^{i+1}}$ and calculate the shape cost volume ${c_{1}^{i}}$ using ${\dot{\gamma}_{2}^{i}}$ and ${\gamma_{1}^{i}}$ as described in [27]. We then estimate the shape flow residual ${\nabla\dot{f}_{S}^{i}}$ using four combinations {Conv2d+LeakyRelu} and add it to ${f_{S}^{i}}$ to obtain the shape-aware flow ${\dot{f}_{S}^{i}}$. At the content-aware stage, we similarly warp ${\delta_{1}^{i}}$ to ${\dot{\delta}_{1}^{i}}$ via ${\dot{f}_{S}^{i}}$ and calculate the content cost volume ${c_{2}^{i}}$ using ${\dot{\delta}_{1}^{i}}$ and ${\gamma_{1}^{i}}$. The content flow residual ${\nabla\ddot{f}_{S}^{i}}$ is estimated using another four combinations {Conv2d+LeakyRelu} and added to ${\dot{f}_{S}^{i}}$ to obtain the upsampled output ${f_{S}^{i}}$. Moreover, the first SCAW module initializes the flow by calculating ${c_{1}^{i}}$ and ${c_{2}^{i}}$ from ${\left\{{\gamma_{1}^{i},\gamma_{2}^{i}}\right\}}$ and ${\left\{{\gamma_{1}^{i},\delta_{1}^{i}}\right\}}$.
@@ 3.3.4 Objective Function
${{\ell_{wi}}}$ supervises the global shape matching in the initial aligning phase, consists of two items, and is defined as (5). The former item punishes the shape error between ¯${\bar{\cal M}_{B}^{hc}}$ and ${{\cal M}_{SAB}^{hc}}$, whereas the latter encourages the garment pixel ratio of ¯${\bar{\cal M}_{B}^{hc}}$ to be similar to ${{\cal M}_{SAB}^{hc}}$, where ${H}$ and ${W}$ represent the height and width of masks; ${{\lambda_{4}}}$ and ${{\lambda_{5}}}$ are weights.
@@
${{\ell_{wr}}}$ is the penalty term of the flow displacement, which consists of two items as (7). We exploit both the one-order constraint in [11] and the second-order constraint in [9] to be ${{\ell_{wr1}}}$ and ${{\ell_{wr2}}}$. ${{\lambda_{9}}}$ and ${{\lambda_{10}}}$ are weights.
@@ 3.4 Arm Regrowth
The upper body skin category has two cases in garment transfer: one case is to transfer a short-sleeve garment onto a person who wears a long-sleeve garment, and another case is to transfer a long-sleeve garment onto a person who wears a short-sleeve garment. To distinguish these two cases precisely, we compare the source and target shapes of upper body skin category, i.e. ${{\cal M}_{A}^{ha}}$ and ${{\cal M}_{SAB}^{ha}}$. The former case is defined as ${{\cal M}_{A}^{ha}>\left({{\cal M}_{SAB}^{ha}\cup{\cal M}_{A}^{ha}}\right)}$ while the latter case is defined as ${{\cal M}_{A}^{ha}<\left({{\cal M}_{SAB}^{ha}\cup{\cal M}_{A}^{ha}}\right)}$. For the former case, we just inherit the part of existing skin. For the latter case, we need to design a model to infer new exposure skin, and name this task ‘arm regrowth’.
Figure 7: The training phase of ${{G_{r}}}$.
As shown in Fig. 7, we train the arm regrowth model ${{G_{r}}}$ in a self-supervision manner, where ${{G_{r}}}$ is implemented by a ResUnet [29]. In the training phase, we adjust both two cases into an arm regrowth task with a ground truth. For the former case, we swap inputs, take ${{\cal M}_{SAB}^{ha}\odot{\cal I}_{A}^{ha}}$ and ${{\cal M}_{A}^{ha}}$ as the source content and target shape, and leverage ${{\cal M}_{A}^{ha}\odot{\cal I}_{A}}$ as the ground truth. ${{G_{r}}}$ predicts ~${\tilde{\cal I}_{A}^{ha}}$ to simulate the arm regrowth task as ~${\tilde{\cal I}_{A}^{ha}={G_{r}}\left({{\cal M}_{SAB}^{ha}\odot{\cal I}_{A}^{ha},{\cal M}_{A}^{ha}}\right)}$. For the latter case, though inputs conform to the arm regrowth task, the ground truth is lacking. Therefore, we randomly crop ${{\cal M}_{A}^{ha}}$ to ˙${\dot{\cal M}_{A}^{ha}}$, take ˙${\dot{\cal M}_{A}^{ha}\odot{{\cal I}_{A}}}$ and ${{\cal M}_{A}^{ha}}$ as the source content and target shape, and utilize ${{\cal M}_{A}^{ha}\odot{\cal I}_{A}}$ as the ground truth. Then, ${{G_{r}}}$ predicts ~${\tilde{\cal I}_{A}^{ha}}$ as ~${\tilde{\cal I}_{A}^{ha}={G_{r}}\left({\dot{\cal M}_{A}^{ha}\odot{{\cal I}_{A}},{\cal M}_{A}^{ha}}\right)}$. Moreover, the cropping strategy is to remove half-plane shape on the horizontal or vertical axis, and these horizontal or vertical probability is equal.
In the testing phase, for the former case, we multiply ${{\cal M}_{SAB}^{ha}}$ with ${{\cal M}_{A}^{ha}\odot{{\cal I}_{A}}}$ to form ${{\cal I}_{SAB}^{ha}}$, and for the latter case, we use pre-trained ${{G_{r}}}$ to predict ${{\cal I}_{SAB}^{ha}}$ via ~${\tilde{\cal I}_{SAB}^{ha}={G_{r}}\left({{\cal I}_{A}^{ha},{\cal M}_{SAB}^{ha}}\right)}$.